# Floquet breathers in a modulated nonlinear lattice

### v2

@@ Abstract
In this work, we study a space-time modulated electro-mechanical system, consisting of an array of coupled cantilevers with their on-site potential provided by electromagnets driven by AC currents. Model equations are derived, and the effect of the modulation on the dispersion bands is examined. The theory of breather existence and stability is extended to include space-time modulation. We perform numerical simulations in a time-modulated system, showing three types of breather response depending on the driving frequency: (i) the modulation frequency is an integer multiple of the breather frequency or, in other words, this phenomenon corresponds to period doubling, tripling, etc.; (ii) the opposite, that is, the breather frequency is an integer multiple of the modulation frequency, corresponding to period-halving, etc. (iii) the breather and modulation frequencies are commensurate in a different form. We use for all of them the term Floquet breathers in analogy with Floquet solitons in photonic systems. As there is no dissipation, but periodic forcing, the energy is conserved but only at discrete times. There exists in this system a huge variety of breathers, either site-centered, symmetric and antisymmetric, bond centered, in-phase or in-quadrature with the modulation, and we analyze the evolution of stability of some of them as a function of the modulation frequency. The construction of a similar system would be of interest to study the properties of dynamic metamaterials.
@@ I Introduction
Highly localized periodic vibrations, so-called intrinsically localized modes (ILM) or discrete breathers (DB), are ubiquitous in many nonlinear lattice-type systems. First reported in  Sievers and Takeno (1988), ILMs have been shown to exist with different spatial patterns; the simplest one is the Sievers-Takeno (ST) mode, where a single oscillator, called the center, has an amplitude much larger than its nearest neighbors, so the amplitude decreases rapidly with the distance to the center. Other mode is the Page (P) mode Page (1990); Flach and Gorbach (2005), where two neighboring oscillators have the same amplitude, either in-phase or anti-phase, being therefore a double ILM or breather.
An important issue regarding ILMs concerns their stability. In Fermi-Pasta-Ulam-Tsingou (FPUT) lattices with an anharmonic interaction, ST modes are unstable, while P modes are often stable. The stability of these modes is related to their spatial symmetry and the Peierls-Nabarro potential barrier Flach and Gorbach (2005). In contrast, the stability properties of these modes depend on the class of lattice model considered. In lattices with nonlinear on-site potentials, such as nonlinear Klein-Gordon (NKG) systems, rigorous results based on the anti-continuum limit show that ST modes exist and are linearly stable under appropriate non-resonance conditions, irrespective of whether the on-site potential is soft or hard, where softness (hardness) corresponds to a decrease (increase) of the oscillation frequency with increasing amplitude. The stability of P modes in such systems is more subtle and may involve stability exchange with ST modes as system parameters are varied Flach and Gorbach (2008).
@@
The study of breather stability were considerably enhanced by Aubry’s band theory, where the eigenvalues of the Newton operator, that is, the operator corresponding to the perturbation of a breather, allow for precise determination of the stability and structural stability of breathers and multibreathers Aubry (1997).
@@
In this work, we are interested in breathers in a time-periodic nonlinear system. We will use the term Floquet breathers, similar to Floquet solitons used in photonic systems Mukherjee and Rechtsman (2020, 2023); Kang et al. (2025); Parker et al. (2022); Johansson et al. (2025). Therefore, Floquet breathers are time-periodic, spatially localized vibrations in discrete, time-periodic nonlinear systems. The breather period does not need to be the period of the system, but commensurate with it as used in the last references. We will consider a specific time-periodic nonlinear system as described below.
@@
Arrays of coupled cantilevers have played a relevant role in the study of nonlinear vibrations, in particular on energy localization Sato et al. (2006), and are among the physical systems where breathers have been reported. They can be built at very different sizes, including nano-, micro- and macro-scales, and can be driven by different types of forces. They are part of important applications, such as actuators and sensors. An experimental macroscopic array of cantilevers was presented in Ref. Kimura and Hikihara (2009a). The system was driven by a periodic force applied to the support of the cantilevers, appearing as an independent external (additive) force in the dynamical equations. The restoring force had two contributions: one provided by the elastic properties of the cantilever, and other of magnetic nature, by fixing a magnet at the free end of the cantilever, and locating an electromagnet below it. The magnetic force due to permanent magnet-electromagnet interaction, can be tuned by changing an electric DC current, in this way, a tunable on-site potential is obtained.
@@
The system resulted in breathers having different properties depending on the current magnitude and the frequency. These and similar models have been used in the study of ILMs Sato and Sievers (2007); Kimura and Hikihara (2009b); Kimura et al. (2016); Chong et al. (2019); Araki and Hikihara (2024).
@@
In this paper we present a variation of the model introduced in Ref. Kimura and Hikihara (2009a), but incorporating, besides the static (DC) current, an oscillating (AC) contribution in the electromagnets. This brings about an on-site potential that it is tunable and periodic in time. Furthermore, introducing phase differences between neighboring cantilevers, an on-site potential with a space-time modulation is achieved. An experimental and mathematical model with a linear on-site potential and nonlinear coupling has been studied Chong et al. (2024), finding $q$-gap breathers. In our case the on-site potential is nonlinear and the coupling is linear.
@@
The article is organized as follows: after the introduction in Sec. I, Sec. II presents the physical model and the derivation of the system of differential equations, both for the nonlinear system and its linear approximation corresponding to small oscillations. In Sec. III, the effect of space and time modulation on the dispersion bands is presented, while the theoretical deduction is located in Appendix B. Sec. IV presents changes brought about by modulation in breather theory, breather obtention, and stability. Sec. V, presents breathers with a period that is an integer of the modulation period, and thus with a frequency smaller than the modulating one, and their properties are analyzed. The case of the breather frequency as a multiple of the modulating frequency is presented in Sec. VI. Sect. VII considers also rational ratios of the breather and modulation frequencies and the possibility of the breathers to be in quadrature with the modulation. A systematic study of the evolution of stability with the frequency is performed. The final Sec. VIII presents the concluding remarks. As already stated, the mathematical deduction of the dispersion relations is included in Appendix B, preceded by Appendix A introducing the notation used and some concepts perhaps not familiar to some readers.
@@ II Physical system and dynamical equations
In this section we derive from physical laws the dynamical equations. A sketch of the physical model is shown in Fig. 1.
Figure 1: System of cantilevers mechanically coupled by a rod. (b) Side view of a cantilever. A permanent magnet is attached to the tip of the cantilever. An electromagnet is placed below the permanent magnet. (c) Definition of magnetic charges. Modified with permission from Ref. Kimura and Hikihara (2009a)
The variable $u_{n}$ measures the horizontal deviation of the cantilever tip with respect to its equilibrium position. The vertical distance between the electromagnet and the cantilever tip is $y_{n}$, therefore, the position vector of the tip with respect to the electromagnet is $\vec{r}_{n}=u_{n}\hat{e}_{x}+y_{n}\hat{e}_{y}$, with equilibrium position $\vec{r}_{n,0}=d_{0}\hat{e}_{y}$, and modulus $r_{n}=\sqrt{u_{n}^{2}+y_{n}^{2}}$. Within the formulation of magnetic charges, analogous to the Coulomb law, the magnetic force is given by $\vec{\cal F}(\vec{r}_{n})=\frac{\displaystyle 1}{\displaystyle 4\pi\mu_{0}}\frac{\displaystyle m_{p}m_{e}}{\displaystyle r_{n}^{2}}\hat{r}_{n}$Griffiths (2024) with $m_{p}$ and $m_{e}$ the magnetic charges of the permanent magnet and the electromagnet, respectively, and $\hat{r}_{n}$ the unitary vector in the direction of $\vec{r}_{n}$. Assuming small cantilever angles, $y_{n}\simeq d_{0}$ and the horizontal force component per unit mass becomes
@@
The strength of the magnetic force is controlled by the current $I_{EM}$ of the electromagnet. The first factor in Eq.(1) can be identified with an interaction coefficient (force per unit mass), $\chi(I_{EM})$, which may be alternatively expressed as a function of the driving current $I_{EM}$ as $\chi(I_{EM})=\chi_{0}+\chi_{1}I_{EM}$. The constant term $\chi_{0}$ in the interaction coefficient is due to the ferromagnetic core of the electromagnet, while the variable term $\chi_{1}$ is due to the magnetic field created by the current of the electromagnet. Note that $\chi_{0}<0$, corresponding to an attractive force, and $\chi_{1}$ depends on the polarity of the current, which is such that in this work $\chi_{1}<0$ too. The current consists of a DC and a modulated AC components, i.e.,
@@
Therefore the dynamical equations become:
@@
Typical values of the parameters for the cantilever system, obtained in a previous experiment, are $\omega_{0,0}=2\pi\times 35.72$ rad/s, $d_{0}=3$ $m$, $\chi_{0}=-4.71\times 10^{-5}$ m3s-2, $\chi_{1}=-9.14\times 10^{-3}$m3s-2A-1 and $C=284$ s-2.
@@
$\frac{\displaystyle u_{n}}{\displaystyle(u_{n}^{2}+d_{0}^{2})^{3/2}}=\frac{\displaystyle u_{n}}{\displaystyle d_{0}^{3}}-\frac{\displaystyle 3}{\displaystyle 2}\frac{\displaystyle u_{n}^{3}}{\displaystyle d_{0}^{5}}+o(u_{n}^{5})$ (5)
@@
$-\omega_{0}^{2}\big(-\frac{\displaystyle 1}{\displaystyle{\bar{d}}_{0}^{3}}u_{n}+(1+\delta{\bar{d}}_{0}^{3}\cos(hn-\Omega t))\frac{\displaystyle u_{n}}{\displaystyle(u_{n}^{2}+{\bar{d}}_{0}^{2})^{3/2}}\big)$
@@
for $\bar{d}_{0}=d_{0}/\sigma$. The second term multiplied by $\omega_{0}^{2}=1$ is written in that way because it is completely nonlinear when expanding the terms in $u_{n}$. It also highlights that the linear frequency of the unmodulated system of the isolated oscillator is $\omega_{0}=1$.
@@
$-\omega_{0}^{2}\big(-\delta_{1}u_{n}+(1+\delta_{2}\cos(hn-\Omega t))\frac{\displaystyle u_{n}}{\displaystyle(u_{n}^{2}+{\bar{d}}_{0}^{2})^{3/2}}\big)$
@@
Figure 2: Numerical simulations after thermalization of the linear system. See App. B.1 for details. (Top:) Energy density plot (Bottom) Contour plot of the two-dimensional XTFFT of the coordinates $u_{n}(t)=u(n,t)$ as a function of $n$ and $t$. Parameters: $\kappa=0.0052$, $\delta=0.075$, $\Omega=0.02$, $h=0.5$, $N=128$.
For a representative set of parameters of currents, we fix DC and AC amplitudes as $I_{AC}=I_{DC}=12$ mA, and in order to obtain that $\omega_{0}=1$, the scaling factors are $\tau=4.286$ ms and $\sigma=1.423$ mm, leading to $\bar{d}_{0}=2.1087$, $\kappa=0.0051632$, $\delta_{1}=0.10665=$, $\delta_{2}=0.6996$.

### v3

@@ Abstract
In this work, we study a space-time modulated electro-mechanical system, consisting of an array of coupled cantilevers with their on-site potential provided by electromagnets driven by AC currents. Model equations are derived, and the effect of the modulation on the dispersion bands is examined. The theory of breather existence and stability is extended to include space-time modulation. We perform numerical simulations in a time-modulated system, showing three types of breather response depending on the driving frequency: (i) the modulation frequency is an integer multiple of the breather frequency or, in other words, this phenomenon corresponds to period doubling, tripling, etc.; (ii) the opposite, that is, the breather frequency is an integer multiple of the modulation frequency, corresponding to period-halving, etc. (iii) the breather and modulation frequencies are commensurate in a different form. We use for all of them the term Floquet breathers in analogy with Floquet solitons in photonic systems. As there is no dissipation, but periodic forcing, the energy is generally conserved but only at discrete times. There exists in this system a huge variety of breathers, either site-centered, symmetric and antisymmetric, bond centered, in-phase or in-quadrature with the modulation, and we analyze the evolution of stability of some of them as a function of the modulation frequency. The construction of a similar system would be of interest to study the properties of dynamic metamaterials.
@@
We propose a space-time modulated system formed by an array of cantilevers where the on-site potential is provided by electromagnets fed with DC and AC currents. The modulation changes the phonon bands and the theory of breathers, or localized periodic solutions. We obtain breathers for many different frequencies, larger or smaller than the modulating one, but always commensurate. We analyze in particular the stability dependence on the frequency for equal-period, doubled.period and halved-period breathers. The results are of application for dynamic metamaterials.
@@ I Introduction
Highly localized periodic vibrations, so-called intrinsic localized modes (ILM) or discrete breathers (DB) were first reported by Sievers and Takeno.Sievers and Takeno (1988) They are ubiquitous in many nonlinear lattice-type systems.Flach and Gorbach (2005, 2008) ILMs have been shown to exist with different spatial patterns; the simplest one is the Sievers-Takeno (ST) mode,Sievers and Takeno (1988) where a single oscillator, called the center, has an amplitude much larger than its nearest neighbors, so the amplitude decreases rapidly with the distance to the center. Other mode is the Page (P) mode,Page (1990) where two neighboring oscillators have the same amplitude, either in-phase or anti-phase, being therefore a double ILM or breather.
An important issue regarding ILMs concerns their stability. The stability of ILMs is related to their spatial symmetry and the Peierls-Nabarro potential barrier.Kivshar and Campbell (1993) The stability properties of these modes depend on the class of lattice model considered. In lattices with nonlinear on-site potentials, such as nonlinear Klein-Gordon (NKG) systems, rigorous results on existence and stability of DB has been given.MacKay and Aubry (1994)
@@
The study of breather stability was considerably enhanced by Aubry’s band theory, where the eigenvalues of the Newton operator, that is, the operator corresponding to the perturbation of a breather, allow for precise determination of the linear stability and structural stability of breathers and multibreathers.Aubry (1997)
@@
In this work, we are interested in breathers in a time-periodic nonlinear system. We will use the term Floquet breathers, similar to Floquet solitons used in photonic systems.Mukherjee and Rechtsman (2020, 2023); Kang et al. (2025); Parker et al. (2022); Johansson et al. (2025) Therefore, Floquet breathers are time-periodic, spatially localized vibrations in discrete, time-periodic nonlinear systems. The breather period does not need to be the period of the system, but commensurate with it, as used in the last references. We will consider a specific time-periodic nonlinear system as described below.
@@
Arrays of coupled cantilevers have played a relevant role in the study of nonlinear vibrations, in particular on energy localization,Sato et al. (2006) and they are among the physical systems where breathers have been reported. They can be built at very different sizes, including nano-, micro- and macro-scales, and can be driven by different types of forces. They are part of important applications, such as actuators and sensors.
An experimental macroscopic array of cantilevers was introduced by Kimura and Hikihara in Ref. Kimura and Hikihara, 2009a.The system was driven by a periodic force applied to the support of the cantilevers, appearing as an independent external (additive) force in the dynamical equations. The restoring force had two contributions: one provided by the elastic properties of the cantilever, and other of magnetic nature, by fixing a magnet at the free end of the cantilever, and locating an electromagnet below it. The magnetic force due to permanent magnet-electromagnet interaction, can be tuned by changing an electric DC current, in this way, a tunable on-site potential was obtained.
@@
The system resulted in breathers having different properties depending on the current magnitude and the frequency. These and similar models have been used in the study of ILMs.Sato and Sievers (2007); Kimura and Hikihara (2009b); Kimura et al. (2016); Chong et al. (2019); Araki and Hikihara (2024)
@@
In this paper we present a variation of the model proposed in Ref. Kimura and Hikihara, 2009a, incorporating, besides the static (DC) current, an oscillating (AC) contribution in the electromagnets. This brings about an on-site potential that it is tunable and periodic in time. Furthermore, introducing phase differences between neighboring cantilevers, an on-site potential with a space-time modulation is achieved. An experimental and mathematical model with a linear on-site potential and nonlinear coupling has been studied,Chong et al. (2024) finding $q$-gap breathers. In our case the on-site potential is nonlinear and the coupling is linear.
@@
The article is organized as follows: after the introduction in Sec. I, Sec. II presents the physical model and the derivation of the system of differential equations, both for the nonlinear system and its linear approximation corresponding to small oscillations. In Sec. III, the effect of space and time modulation on the dispersion bands is presented, while the theoretical deduction is located in Appendix C. Sec. IV presents changes brought about by modulation in breather theory, breather obtention, and linear stability. Sec. V, presents breathers with a period that is an integer of the modulation period, and thus with a frequency smaller than the modulating one, and their properties are analyzed. The case of the breather frequency as a multiple of the modulation frequency is presented in Sec. VI. Sec. VII considers also rational ratios of the breather and modulation frequencies and the possibility of the breathers to be in quadrature with the modulation. A systematic study of the evolution of linear stability with the frequency is performed. The final Sec. VIII presents the concluding remarks. Important but long explanations and mathematical derivations are included in three appendices in order not to interrupt the flow of the article focused on Floquet breathers. Appendix A deals with the concept of thermalization and numerical experiments that are used to check the theory. Appendix B present concepts and notation used in some of the main text and in Appendix C, where the mathematical derivations of the dispersion relations are included.
@@ II Physical system and dynamical equations
In this section we derive from physical laws the system dynamical equations. A sketch of the physical model is shown in Fig. 1.
Figure 1: (a) System of cantilevers mechanically coupled by a rod. (b) Side view of a cantilever. A permanent magnet is attached to the tip of the cantilever. An electromagnet is placed below the permanent magnet. (c) Definition of magnetic charges. Modified with permission from Ref. Kimura and Hikihara, 2009a
The variable $u_{n}$ measures the horizontal deviation of the cantilever tip with respect to its equilibrium position. The vertical distance between the electromagnet and the cantilever tip is $y_{n}$, therefore, the position vector of the tip with respect to the electromagnet is $\vec{r}_{n}=u_{n}\hat{e}_{x}+y_{n}\hat{e}_{y}$, with equilibrium position $\vec{r}_{n,0}=d_{0}\hat{e}_{y}$, and modulus $r_{n}=\sqrt{u_{n}^{2}+y_{n}^{2}}$. Within the formulation of magnetic charges, analogous to the Coulomb law, the magnetic force is given by $\vec{\cal F}(\vec{r}_{n})=\frac{\displaystyle 1}{\displaystyle 4\pi\mu_{0}}\frac{\displaystyle m_{p}m_{e}}{\displaystyle r_{n}^{2}}\hat{r}_{n}$,Griffiths (2024) with $m_{p}$ and $m_{e}$ the magnetic charges of the permanent magnet and the electromagnet, respectively, and $\hat{r}_{n}$ the unitary vector in the direction of $\vec{r}_{n}$. Assuming small cantilever angles, $y_{n}\simeq d_{0}$, the horizontal force component per unit mass becomes
@@
The strength of the magnetic force is controlled by the current $I_{EM}$ of the electromagnet. The first factor in Eq.(1) can be identified with an interaction coefficient (force per unit mass), $\chi(I_{EM})$, which may be alternatively expressed as a function of the driving current $I_{EM}$ as $\chi(I_{EM})=\chi_{0}+\chi_{1}I_{EM}$. The constant term $\chi_{0}$ in the interaction coefficient is due to the ferromagnetic core of the electromagnet, while the variable term $\chi_{1}$ is due to the magnetic field created by the current of the electromagnet. Note that $\chi_{0}<0$, corresponding to an attractive force, and $\chi_{1}$ depends on the polarity of the current, which is such that in this work $\chi_{1}<0$ too. The current consists of DC and AC (modulated) components, i.e.,
@@
Therefore the dynamical equations for an array of coupled cantilevers become:
@@
Typical values of the parameters, obtained in a previous experiment, are $\omega_{0,0}=2\pi\times 35.72$ rad/s, $d_{0}=3$ $m$, $\chi_{0}=-4.71\times 10^{-5}$ m3s-2, $\chi_{1}=-9.14\times 10^{-3}$m3s-2A-1 and $C=284$ s-2.
@@
$\frac{\displaystyle u_{n}}{\displaystyle(u_{n}^{2}+d_{0}^{2})^{3/2}}=\frac{\displaystyle u_{n}}{\displaystyle d_{0}^{3}}-\frac{\displaystyle 3}{\displaystyle 2}\frac{\displaystyle u_{n}^{3}}{\displaystyle d_{0}^{5}}+o(u_{n}^{5})\,$, (5)
@@
$-\big(-\frac{\displaystyle 1}{\displaystyle{\bar{d}}_{0}^{3}}u_{n}+(1+\delta{\bar{d}}_{0}^{3}\cos(hn-\Omega t))\frac{\displaystyle u_{n}}{\displaystyle(u_{n}^{2}+{\bar{d}}_{0}^{2})^{3/2}}\big)$
@@
for $\bar{d}_{0}=d_{0}/\sigma$. The second term in parentheses is written in that way because it is completely nonlinear when expanding the terms in $u_{n}$. It also highlights that the linear frequency of the unmodulated system of the isolated oscillator is $\omega_{0}=1$.
@@
$-\big(-\delta_{1}u_{n}+(1+\delta_{2}\cos(hn-\Omega t))\frac{\displaystyle u_{n}}{\displaystyle(u_{n}^{2}+{\bar{d}}_{0}^{2})^{3/2}}\big)$
@@
For a representative set of parameters, we fix DC and AC amplitudes at $I_{AC}=I_{DC}=12$ mA, and in order to obtain that $\omega_{0}=1$, the scaling factors are $\tau=4.286$ ms and $\sigma=1.423$ mm, leading to $\bar{d}_{0}=2.1087$, $\kappa=0.0051632$, $\delta_{1}=0.10665=$, $\delta_{2}=0.6996$.
@@
The corresponding Hamiltonian is given by:
$H=\sum_{n=1}^{N}e_{n}\quad\mathrm{with}$
$e_{n}=\frac{\displaystyle 1}{\displaystyle 2}p_{n}^{2}+V_{n}(u_{n},t)+\frac{1}{2}U(u_{n+1}-u_{n})+\frac{1}{2}U(u_{n}-u_{n-1})\,$, (10)
where $e_{n}$ is the local energy at site $n$, $V_{n}$ the site and time dependent on-site potential, and $U(u_{n+1}-u_{n})$ is the interaction energy between the variables $u_{n+1}$ and $u_{n}$. We consider periodic boundary conditions so as $u_{N+1}=u_{1}$ and $u_{0}=u_{N}$. The on-site potential $V_{n}$ is given by:
$V_{n}(u_{n},t)=$ $\frac{\displaystyle 1}{\displaystyle 2}\omega_{0}^{2}u_{n}^{2}-\frac{\displaystyle 1}{\displaystyle 2}\delta_{1}u_{n}^{2}$
$+$ $(1+\delta_{2}\cos(hn-\Omega t))\left(\frac{1}{d_{0}}-\frac{1}{\sqrt{u_{n}^{2}+d_{0}^{2}}}\right)\,.$ (11)
The interaction potential is quadratic, given by:
$U(u_{n+1}-u_{n})=\frac{1}{2}\kappa(u_{n+1}-u_{n})^{2}\,.$ (12)
For small oscillations the on-site potential becomes also quadratic:
$V_{n}^{L}(u_{n},t)=\frac{\displaystyle 1}{\displaystyle 2}\omega_{0}^{2}u_{n}^{2}+\frac{\displaystyle 1}{\displaystyle 2}\delta\cos(hn-\Omega t))u_{n}^{2}\,.$ (13)
The Hamiltonian equations $\dot{u}_{n}=-\frac{\displaystyle H}{\displaystyle\partial p_{n}}=p_{n}$ and $\dot{p}_{n}=-\frac{\displaystyle\partial H}{\displaystyle\partial u_{n}}$ are still valid and therefore the time derivative of the Hamiltonian becomes:
$\frac{\displaystyle\textrm{d}H}{\displaystyle\textrm{d}t}=\frac{\displaystyle\partial H}{\displaystyle\partial t}=\delta_{2}\Omega\sum_{n}\sin(hn-\Omega t)\Big(\frac{\displaystyle 1}{\displaystyle d_{0}}-\frac{\displaystyle 1}{\displaystyle\sqrt{d_{0}^{2}+u_{n}^{2}}}\Big)\,.$ (14)
Figure 2: Numerical simulations after thermalization of the linear system. See App. C.1 for details. ( Top) Energy density plot ( Bottom) Contour plot of the two-dimensional XTDFT of the coordinates $u_{n}(t)=u(n,t)$ as a function of $n$ and $t$. Parameters: $\kappa=0.0052$, $\delta=0.075$, $\Omega=0.02$, $h=0.5$, $N=128$.
In general this derivative will not be zero and also it will not be periodic in time, implying that $H$ will not be periodic either. For $h=0$, the coefficients of $\sin(\Omega t)$ have all the same negative sign, and $\textrm{d}H/\textrm{d}t=0$ only if all $u_{n}=0$. However $H$ can be conserved at periodic times. If $T_{\text{m}}=2\pi/\Omega$ is the modulation period and the coordinates $u_{n}$ are periodic with frequency $\omega_{\text{b}}$ and period $T_{\text{b}}=2\pi/\omega_{\text{b}}$, a sufficient condition for $H$ to be periodic with period $T_{\text{H}}$ is that there are integers $m_{\text{m}}$ and $m_{\text{b}}$ such that $T_{\text{H}}=m_{\text{m}}T_{\text{m}}=m_{\text{b}}T_{\text{b}}$ or, in terms of the frequencies:
$\frac{\displaystyle\omega_{\text{b}}}{\displaystyle\Omega}=\frac{\displaystyle m_{\text{b}}}{\displaystyle m_{\text{m}}}\,$, (15)
That is, the frequencies $\omega_{\text{b}}$ and $\Omega$ are commensurate. This is the case of stationary breathers and other periodic solutions.
Note that for small oscillation Eq. (14) becomes:
$\frac{\displaystyle\textrm{d}H}{\displaystyle\textrm{d}t}=\frac{\displaystyle\partial H}{\displaystyle\partial t}=\delta\Omega\sum_{n}\sin(hn-\Omega t)\frac{1}{2}u_{n}^{2}\,.$ (16)